A compound with molecular mass 180 is acylated with CH3COCl to get a compound with molecular mass 390. The number of amino groups present per molecule of the former compound is :
A) 2
B) 5
C) 4
D) 6

Grade:12

1 Answers

Naveen Kumar
9 years ago
The acyl group =CH3CO
its molecular weight=12*2+3*1+16=43
The increase in molecular mass of the product=390-180=210
Let n is the number of amino group present in the molecule and so during the reaction, it would get acylated.
So 1 H for each amono group would be displaced for each of the acyl group
so net incease in the mass due to addition of one group=43-1=42
So for n number of groups, 42*n=210
n=5